OpenStudy (anonymous):

MEDALS/FAN Write an equation for the line that is parallel to the given line and passes through the given point. y = x + 8; (2, 16) A. y = x – 6 B. y = x + 6 C. y = x D. y = x + 14

OpenStudy (anonymous):

D

OpenStudy (***[isuru]***):

hi.. since the 2 lines r parallel their slopes r equal. which means the function of the line the question asks have the same slope as in y = x + 8 the slope of y = x + 8 is 1 so the slope of the line which is parallel to it also 1 then the function is something like this... y = x + c now use the coordinates given (2 , 16) and plug those values in to the function.... y = x + c 16 = 2 + c c = 14 so... the line parallel to the given function and passes through the mentioned point is y = x + 14 hope this will help ya!!

OpenStudy (anonymous):

Thank you both! Here's another... Does the equation represent a direct variation? If so, find the constant of variation. 5x = -3y A. yes; k = 5/3 B. no C. yes; k = 3/5 D. yes; k = -5/3

OpenStudy (***[isuru]***):

 I think now u can find the answer....